jhonyy9 (jhonyy9):

The integer n3 + 2n is divisible by 3 for every positive integer n prove it by math induction is it my proof right ?

jhonyy9 (jhonyy9):

n(n^2 +2)/3 for n=1 ---- 1(1+2)=3/3 =1 n=2 ---- 2(4+2)=12/3=4 n=3 ---- 3(9+2)=33/3=11 for n ---- n(n^2 +2) /3 so using mathinduction we write it for n=k and suppos it true what will be k(k^2 +2)/3 suppos it true so than for k=k+1 we get (k+1)((k+1)^2 +2) /3 for k=1 --- (1+1)((1+1)^2 +2) /3 2(2^2 +2) /3 = 2(4+2) /3 =2*6 /3 = 12/3 =4

OpenStudy (campbell_st):

prove it true for n = k which you have done then assume its true for n = k then let \[k(k^2 + 2) = 3p\] which is really saying that the kth them has factors of 3 and some number p prove n = k + 1 so \[(k + 1)((k + 1)^2 + 2) = 3q\] for some integer q then LHS \[(k+1)(k^2 + 2k + 1 +2) \] simplified its \[k^3 +3k^2 + 5k + 3 \] which can be written as \[k^3 + 2k + 3k^2 + 3k + 3 = 3p + 3(k^2 + k + 1)\] now the common factor 3 can be removed so that \[3(p + k^2 + k + 1) = 3q \] therefore left hand side = right hand side.... and \[q = p + k^2 + k + 1\] therefore its true for n = k + 1 hope this helps

jhonyy9 (jhonyy9):

thank you very nice

OpenStudy (anonymous):

we don't put k =1 when proving in math induction

OpenStudy (anonymous):

@aajugdar yes you do. you start with k=1 then assume for k, then check for k+1

OpenStudy (anonymous):

well in here its n=1 n=k and n =k+1 ne I was correcting his equation solving method :D in n=k+1 he did put k=1 again

OpenStudy (anonymous):

@jhonyy9 still confused?

jhonyy9 (jhonyy9):

so thank you for your opinion but this k=1 after the k=k+1 just wann being an example that is true too ,the same for k=1 hope that you will anderstand me thank you bye

OpenStudy (anonymous):

Who needs induction? Let's just take cases! Any positive integer would be divisible by 3, or would leave a remainder of 1 or 2 when divided by 3 :) For any positive integer n, we have... \[\Large n^3+2n=n(n^2+2)\] So, if n is divisible by 3, then clearly, the proof is done :D Suppose n leaves a remainder of 1, when divided by 3. Then there's an integer h, such that... \[\Large n=3h+1\]\[\Large n(n^2+2)=n[(3h+1)^2+2]=n(9h^2+6h+1+2)\]\[\Large =3n(3h^2+2h+1)\]So it's still divisible by 3. Suppose instead, that n leaves a remainder of 2, when divided by 3. Then there's an integer k, such that \[\Large n=3k+2\]\[\Large n(n^2+2)=n[(3k+2)^2+2]=n(9k^2+12k+4+2)\]\[\Large =3n(3k^2+4k+2)\]So it's still divisible by 3. So in all cases, it will be divisible by 3 :) Problem solved :>